From nch@roe.ac.uk Sat Oct 21 11:53:03 2006 Date: Fri, 20 Oct 2006 17:22:19 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , John Taylor , Jim Emerson , Malcolm Stewart , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive project weekly meeting minutes 18/10/06 Minutes of WFAU VDFS Science Archive meeting: 18th October 2006 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, MAR, JB, PMW, BCW Apologies: JPE, JMS, RGM, JDT, NJC, RSC, ETWS, MSH, AL (many folks away at ADASS) DONM: 10am, Friday 27th October 2006 in the Plate Library Actions discharged this week: ----------------------------- p) Release transit survey data as non-survey databases - ACTION: MAR, NCH & JB Discharged and note sent to PIs/AA/AL/SJW. ACTION: PMW and NCH to book their individual travel arrangements in order to arrive Monday 23rd PM for a last minute hymn-sheet check with JPE and CASU in the pub in the evening. Discharged ACTION: PMW, NCH and RSC to book their required individual accomodation in the Arundle House. Discharged ACTION: NCH to forward Helen Xiang's email address to ETWS to set the wheels in motion over an SDSS-DR5 copy. Discharged ACTION: RSC to make cross-neighbour tables between UKIDSS and 2XMMP Discharged; UKIDSS-XMM cross correlations currently being checked by Leicester folks for sanity and usefulness. ACTION: JB to ingest June 05Av3 prior to May in order that UKIDSS can preview. Discharged; in fact all 05A v3 is now in. ACTION: PMW to get advice/quotes on a NAS storage block based on (possibly) bleeding-edge 750 GB disks. Discharged; quotes and options to be discussed next week. Actions partly discharged but continuing: ----------------------------------------- The following from last week partly done but continue: m) rearrange external catalogues onto batch catalogue server - ACTION: JB Actions carried forward from 06/10/06 meeting: ---------------------------------------------- The following from last week continue: d) Debug CU7 for deprecated frame sets and deletions on rerun - ACTION: NCH i) Enhance seaming to use quality bit information - ACTION: RSC k) Ensure ingest code can cope with 06A missing data quirks and any associated new attributes - ACTION: ETWS o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH ACTION: NCH to circulate a release schedule for DR2 to all concerned. - CONTINUES; see Survey Release below. ACTION: NCH to read the CASU overview and pipeline paper documents - CONTINUES ACTION: NJC (on behalf of NCH) to read the calibration document - CONTINUES ACTION: AL to email Peter Clarke in NeSC to see if a little pressure from above helps to expedite implementation of UKLight at the EUCS end. - CONTINUES ACTION: JB and ETWS to scrub the constituent frames of interleaves for archived UKIDSS data only (where it has not been done already). - CONTINUES Specific points and new actions: -------------------------------- Project management: NCH distrubuted the VDFS FDR comments received so far, and the team discussed some responses. ACTION: NCH to forward the review comments to RSC, MAR, RGM and PMW for their perusal. PMW presented the Q4'06 plan; NCH asked for an additional item on enhanced seaming to be included for CU7. PMW will finalise the plan after next week's review for final ratification at next week's meeting. WFCAM update: Magnitude 6.6 earthquake on the Big Island (epicentre worringly near to NCH's holiday hotel for December...): http://tux.wr.usgs.gov/ but AA reports that things are OK with UKIRT and WFCAM is expected to be back on telescope on schedule. Comments and issues arising from CASU fortnightly minutes: The team noted the minutes of the meeting of 9th October 2006, and a small discussion concerning implementation of the astrometric (and photometric) nightly pipeline recalibrations was discussed. MAR reminded NCH that we had previously discussed these and had come up with an outline strategy for dealing with them. Networking: ETWS has been tenaciously attempting the transatlantic DR5 copy based on some new UDP-based stuff supplied by the UIC boys. Surprise surprise, it doesn't work, and they are not amenable to us using the technique we used last time to get DR3. The situation will be progessed when ETWS returns from hols on Monday. MAR raised the question of the SQL Server version under which the transfered backups have been made. ACTION: NCH to email Jan at JHU to enquire about this (and database file sizes at the same time). JB continues to prod the UKLighters, without much response. WSA Operations: 05A v3 all in; UKIDSS heads are perusing for final go/no-go for DR2. Transfer and ingest of 06A continues. Hardware: The fault on the batch catalogue server thutmose was traced to a memory failure; this was fixed very quickly by a combination of Eclipse, JNTD and JB and the server was back online without too much delay (much appreciated by the AstroGrid folks). JB noted that the server room C1 is being powered down for fire system tests this weekend. The archive will be offline from Friday late afternoon until Saturday (on a best-efforts basis) or more likely sometime Monday AM. MAR has put a note on the WSA web site. Software: Not much to report, as most folks are away. Survey Data Release: Nothing to report this week. Non-survey Data Release: MAR noted that the issue of project continuations from semester to semester has arisen; for the time being, we are asking that non-survey PIs register each project for which data is taken so that flat file access can be expedited; ultimately, it is possible to assign data from one project to another for survey-style DB releases but this will have to go into the queue of stuff to do at a low priority. Astrogrid deployment: It was noted that there has been some progress on the secure DSA for authorised access to UKIDSS data through AG and the plan is to have this set up in time for the AG3 PPRP review scheduled for 7th Nov. Miscellaneous: Nothing else this week